A department adds all materials at the beginning of the process and incurs conversion costs uniformly throughout the process. For the month of July, there was no beginning work in process; 39800 units were completed and transferred out; and there were 20000 units in the ending work in process that were 30% complete. During July, $173420 materials costs and $151140 conversion costs were charged to the department.

The unit production costs for materials and conversion costs for July were

Materials Conversion Costs

$3.79 $2.53
$4.36 $3.80
$2.90 $2.53
$2.90 $3.30
Answer #1

Equivalent unit of material = 39800+20000 = 59800

Equivalent unit of conversion = 39800+(20000*30%) = 45800

Cost per equivalent unit of material = 173420/59800 = 2.90

Cost per equivalent unit of conversion = 151140/45800 = 3.30

So answer is d)